A square garden has an area of 841 squared feet, and the gardener wants to install a sprinkler (with a circluar spraying pattern) at the center of the garden. What is the minimum radius of spray the sprinkler would need in order to water all of the garden?

Answer by mananth(16946)

The sprinkling radius will have to be half the length of the diagonal of the square garden.
Garden is a square
The diagonal and the sides of square form a right triangle.
so the diagonal is the hypotenuse with the sides as the other two sides of a right triangle
let the length of side of square be x
area = x*x = x^2
x^2= 841
x= sqrt(841) = 29
side = 29
Apply pythagoras theorem
diagonal ^2= 29^2+29^2
diagonal^2= 1682
diagonal = 41.01
half the diagonal = 20.5 feet
The radius of the sprinkler